OpenStudy (anonymous):

A clock radio is rated as 30 W of power output. If the radio also draws 30 W at 120 V, which will the current draw be? A. 0.25 A B. 4 A C. 3,600 A D. 0.5 A

OpenStudy (anonymous):

power=voltage x current

OpenStudy (anonymous):

30=120*x 30=120x 30/120=120/120x 0.25 = x so it is A